### v1

@@ 6.1.1 Faithfulness
Chain-of-thought and other approaches prompt a language model to take certain steps to solve the problem that the prompt specifies. One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the correct answer (unfaithful reasoning). A few studies measure the faithfulness of reasoning by LLMs. Lanham et al. (2023) notes that just like organic reasoners, a model’s reasoning may be post-hoc, it may be constructed after a certain conclusion has been found. By deliberately adding mistakes to the chain of thought, the authors measure the faithfulness of the model. They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful. Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination.
Another study adds deliberate bias to the prompt. For example, in a multiple-choice setting, they always make answer (a) the correct answer (Turpin et al., 2024). They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers. The authors further note that, insofar as language models are trained on human-written explanations, that explanations may be incomplete or wrong. Human explanations may omit crucial steps of the causal chain, may provide an unfaithful account of the human reasoning process, or may be aimed at convincing others, instead of providing the true causes of a decision.
To address issues of faithfulness, Lyu et al. (2023) propose Faithful-chain-of-thought. This approach involves two stages. First, the natural language query is translated into a formal symbolic language. Second, the problem-solving stage processes the formal language, and can explain the reasoning steps it has thus taken. For the symbolic language, Python, Datalog, or PDDL is suggested. Faithfulness studies tell us more about how models reason. Further surveys on this topic are Mondorf and Plank (2024); Chuang et al. (2024); Luo et al. (2023); Paul et al. (2024),
6.1.2 Scaling
@@
The emergent abilities of LLMs have prompted research into the nature of scaling and reasoning with LLMs, and, specifically, how reasoning capabilities can be transferred to smaller language models. Scaling laws of LLMs are an active area of study, see for example Kaplan et al. (2020); Henighan et al. (2020); Hoffmann et al. (2022). Given the computational cost of LLMs, there is much interest in transferring knowledge to small language models. Comprehensive surveys on knowledge distillation are Xu et al. (2024); Gu et al. (2023). For reasoning specifically, Magister et al. (2022) have studied reasoning in small language models, using a student model that learns from a teacher model, by finetuning. Another study related to Self-taught-reasoner (Li et al., 2022a) focuses on explanation in small language models, achieving similar results.
Other works focus on prompt distillation for retrieval Dai et al. (2022), recommendation (Li et al., 2023), distillation to embodied agents of Chain-of-thought reasoning (Choi et al., ), and distillation of LLM graph reasoning (Zhang et al., 2024). Distillation of reasoning to smaller models can work surprisingly well in situations with more explicit instructions. Distillation is also proposed for bringing results of System 2 reasoning to System 1 Yu et al. (2024), which brings us to the topic of metacognition (see Section 6.2.3).
6.2 Limitations: What LLMs Can and Cannot do
@@
The capabilities of LLMs are impressive. LLMs can be seen as large text-based surrogate models of the world (or the world how we describe it on the internet), and thus allow us to reason in a way that we can understand about a large variety of contexts and problems. Reasoning tasks, such as math word problems, were one of the capabilities that LLMs could not achieve, until recently. Let us look more closely at what language models can and cannot do.
6.2.1 What Can LLMs Do?
@@
With the right prompt LLMs are able to solve many of the problems in reasoning grade school math word benchmarks. Prompt-based learning is able to perform reasoning tasks such as math word problems, robotic movement, and Python code generation, at inference time, without expensive parameter training.
@@
We note that a simple taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature well. Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, or self-verification. Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia. Going a step further, using external control algorithms (such as search or RL) as scaffolding, dynamic prompts can use the LLMs to perform complex and interactive reasoning patterns.
Note that the reasoning control is now two layers away from the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM. This is reasoning with prompts with LLMs, not by.
@@
At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence (AI), symbolic and connectionist.555Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008). Ever since the success of the connectionist approach LeCun et al. (2015); Goodfellow et al. (2016) (deep learning, including LLMs) researchers have tried to join the two approaches. Search and reinforcement learning are rooted in the symbolic AI tradition, while LLMs are rooted in the connectionist tradition. The literature in this survey combines the two traditions. High performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM. In other fields similar combinations can be seen (for example, AlphaFold Bryant et al. (2022); Jumper et al. (2021) and retrosynthesis of molecules Segler et al. (2018)). The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform reasoning. How the two traditions will continue to improve eachother, we will see in further research.
@@
We note that benchmarks such as GSM8K have been central for the progress in the field, and that while reasoning started with math word problems, the field has extended to robotics, autonomous agents, games, and most emphatically computer code. Formal languages play an important role in the intermediate multi-step reasoning chains.
A side effect from the work on reasoning is the emergence of a new few-shot learning approach for sequential decision-making processes (SDP)(Littman, 1996). Traditionally these processes are solved with reinforcement learning (such as DQN Mnih et al. (2015), PPO (Schulman et al., 2017) and SAC Haarnoja et al. (2018)), achieving good results, but suffering from high sample complexity for larger problems Plaat et al. (2023). The emergence of few-shot in-context learning for solving SDPs opens a research avenue to find out what SDPs few-shot prompt-learning will be able to solve.
6.2.2 What Can LLMs Not Do?
Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024). Another line of research argues that LLMs cannot reason, providing examples where LLMs fail, and discussing potential reasons. Berglund et al. (2023) show that LLMs can fail to generalize in surprising ways. They provide the example that if a model is trained to report that ”Valentina Tereshkova was the first woman to travel to space”, it will not automatically be able to answer the question, ”Who was the first woman to travel to space?” pointing to a lack in semantic understanding of LLMs. Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems Wu et al. (2024). The question which problems LLMs can and cannot solve will continue to motivate researchers.
Other works study the dangers of the size of LLMs. Bender et al. (2021) mention the environmental risks associated with the large computational training demands, as well as the difficulty of understanding the training data, for example in the context of bias. Furthermore, there are ethical, legal, and copyright concerns regarding the data that LLMs are trained on. Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not).
@@ 6.2.2 What Can LLMs Not Do?
Most of the reasoning capabilities exhibited by LLMs are due to the great representational powers of the transformer architecture, and how in-context learning is able to harness them. Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers. Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own. Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts. Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm. It is an open question if prompt learning is able get the LLM to create a prompt to exhibit non-trivial reasoning by itself.
@@
From the symbolic planning field there is also a critical view on the reasoning and planning abilities of LLMs (Valmeekam et al., 2023) giving examples of planning failures. They argue that LLMs can be used instead to improve heuristic elements of traditional planners, such as PDDL (Kambhampati et al., 2024), to strengthen traditional symbolic planning approaches.
@@
Some of the names of the approaches surveyed in this paper are suggestive of self-awareness and self-reflective capabilities. True self-reflection, or metacognition, is still largely outside the capabilities of current LLMs. LLMs can be prompted to reason, to take small steps, to self-evaluate, and their search process can be controlled by an external algorithm. The self-reflective type of “intelligence” is written into the prompt by the prompt engineer or the interactive algorithm. We are unaware of any LLM that has been made to reflect on, or even control, its reasoning processes, controlling how many reasoning steps it should take, or limiting its reasoning once the answer had become good enough. True self-reflection remains future work, although some steps have been taken, as we will discuss next.
6.2.3 Reasoning towards Metacognition
@@
Human thought exhibits the ability to reason about self, we are able to think about our own thinking processes. Metacognition studies these topics (Veenman et al., 2006). Prompted by the success of Chain-of-thought and the works that we have surveyed, metacognition has also been studied in the context of LLMs (Toy et al., 2024).
Many reasoning approaches highlight self-reflective aspects in their names and in how they work. The prompts that prompt the models to reason are being improved with the outcome of the reasoning process, and in Buffer-of-thoughts thought-templates are used that are derived from other reasoning processes. Wang and Zhao (2023) study Metacognitive-prompting. Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models. Figure 26 illustrates the relation between metacognitive human thought processes and metacognitive LLM prompting.
@@
Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs. Didolkar et al. (2024) study metacognitive capabilities of LLMs in mathematical problem solving, both on GSM8K and on the harder MATH problems (Hendrycks et al., 2021). First, the model is prompted to find a skill name for each problem instance in the dataset. For 7000 instances of GSM8K, 500 skill names were found by the model. Next, these 500 names are clustered down to 22 skills. They find that by using the names of these 22 skills in Chain-of-thought-like prompts, more problems are solved than with standard Chain-of-Thought/Self-consistency/PAL prompts. Examples of the 22 skill names are multiplication-and-addition, basic-arithmetic, subtraction, and algebra. Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3). The performance of the weak model benefits from the skill-name-enhanced prompts.
The connection between reasoning in LLMs and full-blown metacognitive reasoning is in its early stages. Exciting future research may appear.
6.3 Research Agenda
@@
At the end of this discussion, we present promising topics for future work. Reasoning with LLMs is an active field of research. It brings together elements of symbolic reasoning, connectionism, natural language, autonomous agents, and affective reasoning (Broekens et al., 2023) with the promise of artificial general intelligence.
For the future, the surveyed works point in the following directions. First we discuss topics for the field of LLM-reasoning itself, then we discuss more general machine learning topics that are important for progress in LLM-reasoning, and finally we discuss more longer term, fundamental topics.
Specific research topics for reasoning with LLMs are:
Control and prompt-learning—Search control beyond greedy search is implemented as an external algorithm. Is it possible to incorporate all stages of the reasoning pipeline into an interactive prompt? Can we make a prompt that performs dynamic search-like step control without external scaffolding?
@@
Grounding—Reasoning in LLMs has been successfully applied in autonomous agents, robotics, and games. A challenge is the grounding of the reasoning process in the environment. How can we help LLMs to actively find new information when the reasoning outcome is uncertain? Is retrieval augmented generation the future? Is the future of the reasoning-LLM a search engine (Verberne, 2024)?
@@
Benchmarks—Progress in LLMs is governed by the availability of the right benchmarks. The current favorite is GSM8K, for grade school math. As the field progresses, other benchmarks will become prevalent: benchmarks with more difficult tasks, and benchmarks for other applications in autonomous agents and robotics.
@@
Small language models—Efficiency is an important element for wide adoption of language models. Important topics are distillation of reasoning to small language models and an understanding of scaling laws.
Few-shot Reinforcement Learning—Small reasoning problems can be solved with few-shot in-context learning. Can we solve larger sequential decision processes, reducing the sample complexity in reinforcement learning?
@@
Symbolic and Connectionist Computation—How can we further improve LLM-reasoning: how can LLMs benefit from symbolic reasoning prompts and how can LLMs help ground symbolic reasoning in language?
Metacognition—Much of the research into reasoning guides the model how it should solve a problem. Is it helpful to introduce named concepts for different kinds of reasoning? Can the model find these concepts by itself? Making the LLM “think” step by step is a first step towards influencing the model’s own “thought” processes. The first works on LLM metacognition have appeared, and artificial general intelligence will pursue this further.
7 Conclusion

### v2

@@ 4.2.1. Faithfulness
Chain-of-thought approaches prompt a language model to take certain steps to solve the problem that the prompt specifies. One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the same answer (unfaithful reasoning). A few studies measure the faithfulness of reasoning with LLMs. Lanham et al. (2023) notes that just like organic reasoners, a model’s reasoning may be post-hoc, it may be constructed after a certain conclusion has been found. By deliberately adding mistakes to the chain of thought, the authors measure the faithfulness of the model. They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful. Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination (Huang et al., 2025).
Another study adds deliberate bias to the prompt. For example, in a multiple-choice setting, they always make answer (A) the correct answer (Turpin et al., 2024). They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers. To address issues of faithfulness, Lyu et al. (2023); Xu et al. (2024a) propose Faithful-chain-of-thought. This approach involves two stages. First, the natural language query is translated into a formal symbolic language. Second, the problem-solving stage processes the formal language, and can explain the reasoning steps it has thus taken. For the symbolic language, Python, Datalog, or PDDL is suggested. Another approach, mechanistic interpretability, studies methods to target individual representations inside the LLM, to see if the expected behavior occurs in practice (Rai et al., 2024; Bereska and Gavves, 2024; Chen, 2025).
4.2.2. Scaling
@@
The emergent abilities of LLMs have prompted research into the nature of scaling and reasoning with LLMs, and, specifically, how reasoning capabilities can be transferred to smaller language models. Scaling laws of LLMs are an active area of study, see for example (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). Distillation of reasoning to smaller models can work surprisingly well in situations with more explicit instructions, and given the computational cost of training LLMs, there is much interest in transferring knowledge to small language models. Comprehensive surveys on knowledge distillation are Xu et al. (2024b); Gu et al. (2023). For reasoning specifically, Magister et al. (2023) have studied reasoning in small language models, using a student model that learns from a teacher model, by finetuning. Other works focus on prompt distillation for retrieval (Dai et al., 2023), recommendation (Li et al., 2023a), embodied agents (Choi et al., ), and LLM graph reasoning (Zhang et al., 2024).
Another study related to Self-taught-reasoner focuses on explanation in small language models, also achieving good results of knowledge transfer (Li et al., 2024).
4.3. Limitations: What LLMs Can and Cannot do
@@
The capabilities of LLMs are impressive. LLMs can be seen as large text-based surrogate models of the world (or the world how we describe it on the internet), and thus allow reasoning about a large variety of contexts and problems. Reasoning, such as in math word problems, were one of the capabilities that LLMs could not achieve, until recently. Let us look more closely at what language models currently can and cannot do.
4.3.1. What Can LLMs Do?
@@
With the right prompt, LLMs are able to solve many of the problems in grade school math word benchmarks and beyond. Prompt-based learning is able to perform reasoning tasks such as math word problems, robotic movement gaming, and Python code generation, at inference time, without expensive parameter training.
@@
A taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature. Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, and self-verification. Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia. Going a step further, using external control algorithms as scaffolding (such as search or reinforcement learning), dynamic prompts can use the LLMs to perform complex and dynamic reasoning patterns. Note that the reasoning control is now two layers away from the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM. This is reasoning with prompts with LLMs, not by.
@@
At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence, symbolic and connectionist.444Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008). Ever since the success of the connectionist approach (LeCun et al., 2015; Goodfellow et al., 2016) (deep learning, including LLMs) researchers have tried to join the two approaches. Search and reinforcement learning are rooted in the symbolic tradition, while LLMs are rooted in the connectionist tradition. The literature in this survey combines the two traditions. High performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM. In other fields similar combinations can be seen (for example, AlphaFold (Bryant et al., 2022; Jumper et al., 2021) and retrosynthesis of molecules (Segler et al., 2018)). The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform dynamic reasoning.
@@
We note that benchmarks such as GSM8K have been central for the progress of the field, and that while reasoning started with math word problems, the field has extended to robotics, autonomous agents, games, and most emphatically computer code. Formal languages play an important role in the intermediate multi-step reasoning chains.
4.3.2. What Can LLMs Not Do?
@@
Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024). Most of the reasoning capabilities exhibited by LLMs are due to the representational powers of the transformer architecture, and how in-context learning is able to harness them. Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers. Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need careful evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own. Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts (Plaat et al., 2025). Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm. Studies into in-context reinforcement learning aim to answer the question if prompt learning is able get the LLM to create a prompt to exhibit dynamic reasoning by itself (Lee et al., 2023; Schultz et al., 2024; Demircan et al., 2024; Paglieri et al., 2024).
@@
Some studies on symbolic planning are critical on the abilities of LLMs (Valmeekam et al., 2023), and show examples of planning failures, arguing that LLMs are better used to improve heuristic elements of traditional planners, such as PDDL (Kambhampati et al., 2024), to strengthen traditional symbolic planning approaches.
Other works study the dangers of the size of LLMs. Bender et al. (2021) mention the environmental risks associated with the large computational training demands, as well as the difficulty of understanding the training data, for example in the context of bias. Furthermore, there are ethical, legal, and copyright concerns regarding the data that LLMs are trained on. Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not). Here, mechanistic interpretability can be used to explore LLM representations and understand where they go wrong (Bereska and Gavves, 2024; Rai et al., 2024; Chen, 2025).
@@
Some of the names of the approaches surveyed in this paper are suggestive of self-awareness and self-reflective capabilities. True (human) self-reflection, or metacognition, is still largely outside the capabilities of current LLMs. LLMs can be prompted to reason, to take small steps, to self-evaluate, and their search process can be controlled by an external algorithm. The self-reflective type of “intelligence” is written into the prompt by the prompt engineer or the control algorithm. We are unaware of any LLM that has been made to reflect on, or even control, its reasoning processes, controlling how many reasoning steps it should take, or limiting its reasoning once the answer had become good enough. True self-reflection remains future work, although some steps have been taken, as we will discuss next.
4.3.3. Reasoning towards Metacognition
@@
Human thought exhibits the ability to reason about self, we are able to think about our own thinking processes. Metacognition studies this phenomenon (Veenman et al., 2006). Prompted by the success of Chain-of-thought and related works, metacognition has also been studied in the context of LLMs (Toy et al., 2024).
@@
Many reasoning approaches highlight self-reflective aspects in their names and in how they work. The prompts that prompt the models to reason are being improved with the outcome of the reasoning process, and in Buffer-of-thoughts thought-templates are used that are derived from other reasoning processes. Wang and Zhao (2024) study Metacognitive-prompting. Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models. Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs. Didolkar et al. (2024) study metacognitive capabilities of LLMs in mathematical problem solving, both on GSM8K and on the harder MATH problems (Hendrycks et al., 2021). First, the model is prompted to find a skill name for each problem instance in the dataset. For 7000 instances of GSM8K, 500 skill names were found by the model. Next, these 500 names are clustered down to 22 skills. They find that by using the names of these 22 skills in Chain-of-thought-like prompts, more problems are solved than with standard Chain-of-Thought/Self-consistency/PAL prompts. Examples of the 22 skill names are multiplication-and-addition, basic-arithmetic, subtraction, and algebra. Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3). The performance of the weaker model benefits from the skill-name-enhanced prompts. Begus et al. (2025) propose a program for the study of metalinguistic abilities of LLMs. LLMs struggle with analogies (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990) and analogical reasoning. Metacognitive reasoning with LLMs is still in its early stages.
4.4. LLMs that Generate Computer code
A persistent thread in the results in this survey is that LLMs are good at generating formal languages, such as PDDL, logic, math equations, and Python. The agentic approach in which LLMs are combined with external tools such as interpreters or debuggers often yields good performance (Plaat et al., 2025). LLMs can be used in different modes, which we will now discuss.
When LLMs are used in direct mode, a prompt is provided with an instruction, and an answer. This is the mode in which most casual users use an LLM.
LLMs can also be used in algorithmic mode, or Chain-of-thought mode, where an extra element is added to the prompt that gives an example of which steps to take to arrive at the answer (Sel et al., 2023). Lee et al. (2023); Wang et al. (2024c); Schultz et al. (2024) show that the transformer architecture can be trained such that search and reinforcement learning algorithms can be executed in-context. (LLMs pretrained on natural language are typically not good at in-context reinforcement learning (Ruoss et al., 2024; Paglieri et al., 2024)).
We have seen that LLMs are not good at preserving state between calls (Kempinski et al., 2025). External algorithms can be added to manage the state, that call the LLM in an external mode (Yao et al., 2024, 2023). Tree-of-thoughts, and similar approaches, use the LLM in a stateless functional way, with the prompt containing the instructions, not unlike a natural language variant of a functional programming language such as LISP or Haskell.
Finally, an LLM can be used in code generation mode. The prompt specifies a problem, and the LLM is asked to generate computer code (such as Python or PDDL). The code may contain variables that manage the state correctly. The code can be executed directly, or first be checked and improved by a programmer. We expect that the use LLMs as code generators, and prompt-engineering as natural-language-programming, will continue to grow.
4.5. Research Agenda
@@
At the end of this discussion, we list promising topics for future work. Reasoning with LLMs is an active field of research. It brings together elements of symbolic reasoning, connectionism, natural language, autonomous agents, affective reasoning (Broekens et al., 2023) and metacognition. First we discuss current topics for the field of LLM-reasoning itself, then we discuss more general machine learning topics that are important for progress in LLM-reasoning, and finally we discuss more longer term, fundamental topics.
In Context Reinforcement Learning—Search control beyond greedy search is often implemented as an external reinforcement learning algorithm. Is it possible to incorporate the control stage of the reasoning pipeline into one static prompt, as in-context reinforcement learning (ICRL)? Studies indicate that models can be trained for decision making (Lee et al., 2023; Demircan et al., 2024), but general LLMs have not yet been shown to perform ICRL well (Schultz et al., 2024; Paglieri et al., 2024; Ruoss et al., 2024; van Wetten et al., 2025).
@@
Grounding—Reasoning in LLMs has been successfully applied in autonomous agents, robotics, and games. A challenge is the grounding of the reasoning process in the environment. Retrieval augmented generative methods (RAG) can help LLMs to actively find new information when the reasoning outcome is uncertain. RAG, search, and agentic LLMs, are also active areas of research (Dai et al., 2023; Verberne, 2024; Plaat et al., 2025).
@@
Benchmarks—Progress in LLMs depends on the availability of the right benchmarks. As the field has progressed beyond math word problems, other benchmarks become prevalent, with more difficult and diverse tasks.
@@
Smaller language models—Efficiency is an important element for wide adoption of language models. Smaller models have many advantages over larger models: a smaller environmental footprint, more accessible to people and organizations with fewer computational resources, and the possibility to finetune them. Models may have fewer parameters, or the parameters may be quantized with fewer bits. Unfortunately, with a smaller number of parameters, the models also have less reasoning power. It is important for future work to investigate this trade-off. Solutions to compensate the limited capabilities of smaller models are distillation of reasoning to small language models and reinforcement learning approaches (Shao et al., 2024) to improve the efficiency of the finetuning/reasoning pipeline (Guo et al., 2025; Du et al., 2025).
@@
Symbolic and Connectionist Computation—How can we further improve LLM-reasoning: how can LLMs benefit from reasoning prompts based on the symbolic AI literature, and how can LLMs help ground symbolic reasoning in language?
Multimodal World-Models with Embodied Grounding—How can an LLM maintain a unified, continually updated representation that integrates text, vision, audio, and sensorimotor feedback, enabling it to reason over events across modalities, and refine its world-model through closed perception–action loops in real-world environments?
Norm-Sensitive and Value-Aligned Reasoning—How do we represent, adapt, and audit diverse cultural norms and ethical values within the reasoning process, ensuring that each step of an LLM’s chain-of-thought respects context-dependent moral constraints while remaining transparent and verifiable?
Reasoning in other languages than English—The majority of the research into LLM reasoning is for English. Although there are efforts in creating benchmarks to test LLM capabilities in other languages555https://github.com/NaiveNeuron/awesome-multilingual-llm-benchmarks, these contain mainly NLP tasks such as question answering (QA) and natural language inference (NLI). In addition, challenges related to low-resource languages (languages for which very limited training data is available) have not been addressed yet.
Reasoning in specialized domains—In an effort to guide development and evaluation of new methods, research in AI has a strong focus on benchmarking: standardized datasets with a limited set of problems that are realistic to evaluate. In the real world, however, reasoning problems also occur in more challenging contexts. A few examples are: legal reasoning for the interpretation of case law or writing contracts; scientific reasoning for advancing scientific fields; and medical reasoning for AI-assisted diagnostics.
5. Conclusion